The letters P, Q and R each stands for a different digit in the alphametric (PP)^2 + (PQ)^2 = QRQR. (Here, PP means a two-digit number, like 33.) If PP and PQ are consecutive numbers, then the value of P+Q + R is
Found 3 solutions by math_tutor2020, greenestamps, ikleyn:
Answer by math_tutor2020(3817)


Hint:

1^2 = 1
2^2 = 4
3^2 = 9
4^2 = 16
5^2 = 25
6^2 = 36
7^2 = 49
8^2 = 64
9^2 = 81
Focus on the units digit of each result to get this sequence: 1, 4, 9, 6, 5, 6, 9, 4, 1
Notice we have symmetry.

Sort the values and toss out any duplicates to get: 1, 4, 5, 6, 9

-------------------------------------------------------

Another hint:

Q = P+1 since PP and PQ are consecutive numbers (eg: 64 and 65)
PP and PQ are of different parity (one is odd, the other even, in either order)
(PP)^2 and (PQ)^2 are of different parity. I'll leave the proof for the reader.

odd + even = odd
which means QRQR is odd, and R itself is odd
The possible values of R is from the set {1,3,5,7,9}

Answer by greenestamps(13200)



There are algebraic techniques that you could use to solve the problem, or to work parts of the problem. But the conditions of the problem allow only a very few possible solutions, so trying each of them to find the answer is as good a method as any.

The conditions of the problem say that the number PP is a 2-digit integer with both digits the same, an PQ is the next larger integer. Try each possibility:

11^2+12^2 -- don't even need to do the actual calculations; the sum is not 4 digits
22^2+23^2 = 484+529 = 1013
33^2+34^2 = 1089+1136 = 2225
44^2+45^2 = 1936+2025 = 3961
55^2+56^2 = 3025+3136 = 6161

There is our answer. P = 5; Q = 6; R = 1

ANSWER: P+Q+R = 12


Answer by ikleyn(52783)

.
The letters P, Q and R each stands for a different digit in the alphametric (PP)^2 + (PQ)^2 = QRQR.
(Here, PP means a two-digit number, like 33.) If PP and PQ are consecutive numbers, then the value of P+Q + R is
~~~~~~~~~~~~~~~~~~

First notice that since the sum QRQR is a 4-digit number, the numbers PP and PQ can not be too large.

Indeed, P must be less than 7; otherwise the sum  (PP)^2 + (PQ)^2 will be 5-digit number.



Having it, there are not so many options for you to check: all possible oprtions are

(1)    PP = 66  PQ = 67

(2)    PP = 55  PQ = 56

(3)    PP = 44  PQ = 45

(4)    PP = 33  PQ = 34

(5)    PP = 22  PQ = 23

(6)    PP = 11  PQ = 12


It gives for  (PP)^2 + (PQ)^2  these possible values

(1)                8845

(2)                6161

(3)                3961

(4)                2245

(5)                1013

(6)                 265


Of these options, only (2) has the reqired form, with the number 6161.


So, P = 5, Q = 6, R = 1,

and P + Q + R = 5 + 6 + 1 = 12.    ANSWER

Solved.